Three numbers are selected at random WITHOUT replacement from the first 6 natural numbers. Let X denote the LARGEST of the three numbers obtained. Find the probability distribution of X.

Grade:11

1 Answers

Khimraj
6 years ago
x                  P(x)
1                   0
2                   0
3                  (2C2*1)/6C3 = 1/20
4                  (3C2*1)/6C3 = 3/20
5                  (4C2*1)/6C3= 6/20
6                  (5C2*1)/6C3 = 10/20.
Hope it clears. If you like answer then please approve it.